Use VBA to create a table of FX FX in Excel for 0 <xs 10 incrementing x by 0.1. FX EX
Answer #1

Sub CreateTable()

   

    Dim wsTest As Worksheet

   

    '''Sheet Name

    SheetName = "Table"

    Set wsTest = Nothing

   

    On Error Resume Next

    Set wsTest = ActiveWorkbook.Worksheets(SheetName)

    On Error GoTo 0

    

    If wsTest Is Nothing Then

        Worksheets.Add.Name = SheetName   '''If sheet is not exist

    End If

    Sheets(SheetName).Select

   

    '''Columns Names

    Range("A1").Value = "x"

    Range("B1").Value = "y=x^-2"

    Range("C1").Value = "y=x^-1"

    Range("D1").Value = "y=x^0"

    Range("E1").Value = "y=x^1"

    Range("F1").Value = "y=x^2"

    Range("G1").Value = "y=x^3"

   

    x = 0

    Row = 2

    Do While x <= 10

        If x = 0 Then

            x = x + 0.1

        End If

       

        '''Formulas

        Cells(Row, 1).Value = x

        Cells(Row, 2).Value = x ^ -2

        Cells(Row, 3).Value = x ^ -1

        Cells(Row, 4).Value = x ^ 0

        Cells(Row, 5).Value = x ^ 1

        Cells(Row, 6).Value = x ^ 2

        Cells(Row, 7).Value = x ^ 3

        Row = Row + 1

        x = x + 0.1

    Loop

   

    '''Draw Borders

    Range("A1").Select

    Range(Selection, Selection.End(xlDown)).Select

    Range(Selection, Selection.End(xlToRight)).Select

    Selection.Borders(xlDiagonalDown).LineStyle = xlNone

    Selection.Borders(xlDiagonalUp).LineStyle = xlNone

    With Selection.Borders(xlEdgeLeft)

        .LineStyle = xlContinuous

        .ColorIndex = 0

        .TintAndShade = 0

        .Weight = xlThin

    End With

    With Selection.Borders(xlEdgeTop)

        .LineStyle = xlContinuous

        .ColorIndex = 0

        .TintAndShade = 0

        .Weight = xlThin

    End With

    With Selection.Borders(xlEdgeBottom)

        .LineStyle = xlContinuous

        .ColorIndex = 0

        .TintAndShade = 0

        .Weight = xlThin

    End With

    With Selection.Borders(xlEdgeRight)

        .LineStyle = xlContinuous

        .ColorIndex = 0

        .TintAndShade = 0

        .Weight = xlThin

    End With

    With Selection.Borders(xlInsideVertical)

        .LineStyle = xlContinuous

        .ColorIndex = 0

        .TintAndShade = 0

        .Weight = xlThin

    End With

    With Selection.Borders(xlInsideHorizontal)

        .LineStyle = xlContinuous

        .ColorIndex = 0

        .TintAndShade = 0

        .Weight = xlThin

    End With

   

    '''Heading in Bold

    Range("A1:G1").Select

    Selection.Font.Bold = True

    ActiveWorkbook.Save

    MsgBox ("Done")

   

End Sub 100 10 0.1 0.2 0.1 0.2 0.3 11.11111 3.333333 0.4 0.5 0.4 0.5 0.6 0.7 0.8 0.9 0.01 0.04 0.09 0.16 0.25 0.36 0.49 0.64 0.81 0.001 0.008 0.027 0.064 0.125 0.216 0.343 0.512 0.729 6.25 2.5 4 0.7 2.0408161.428571 0.8 1.5625 1.25 10 12 13 14 15 16 17 1.1 0.8264460.909091 1.2 0.6944440.833333 1.3 0.5917160.769231 1.40.5102040.714286 1.5 0.4444440.666667 0.625 1.71 0.3460211 0.588235 1.8 0.308642 0.555556 1.9 0.2770080.526316 0.5 2.1 0.226757 0.47619 2.2 0.2066120.454545 2.3 0.1890360.434783 2.4 0.173611 0.416667 1.1 1.2 1.3 1.4 1.5 1.6 1.7 1.8 1.9 1.21 1.44 1.69 1.96 2.25 2.56 2.89 3.24 3.61 1.331 1.728 2.197 2.744 3.375 4.096 4.913 5.832 6.859 1.6 0.390625 19 0.25 4 4.41 4.8410.648 5.29 12.167 5.76 13.824 2.1 2.2 9.261 24 25 H 2.4 TableSheet1 Sheet2Sheet3